# Optimization of the Downlink Spectral- and Energy-Efficiency of RIS-aided Multi-user URLLC MIMO Systems

### v1

@@ IV-B2 Impact of the reliability constraint
Fig. 9 shows the average max-min EE versus $\epsilon$ for $P=10$ dB, $N_{{BS}}=4$, $N_{u}=4$, $K=2$, $L=1$, $M=1$, $n_{t}=256$ bits, and $N_{{RIS}}=20$. In this example, the RIS providea a significant gain, which increases with $\epsilon$. Again, RIS decreases the max-min EE when its elements are random. Moreover, we can observe that single-stream data transmission is suboptimal in this $4\times 4$ MIMO system. Indeed, the multi-stream systems communicating without RIS outperforms the RIS-aided single-stream data transmission.
@@ IV-B3 Impact of the packet length
Fig. 12 shows the average EE improvement by RISs versus $n_{t}$ for $P=10$ dB, $N_{{BS}}=2$, $N_{u}=2$, $K=2$, $L=2$, $M=2$, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. The average improvements reduce as $n_{t}$ increases. This indicates that the lower the tolerable latency is, the higher gain the RIS can provide. In other words, RIS-aided systems become more energy efficient when a low latency is required, as in control channels, for example.
V Summary and Conclusions
An optimization framework was proposed for MU-MIMO RIS-aided systems with FBL by considering the NA for the rate expressions. To this end, we first calculated closed-form expressions for the FBL rate and then obtained suitable concave lower bounds for the FBL rates. Our proposed framework can be adapted to any MU-MIMO system in which interference is treated as noise. Moreover, the framework can obtain a stationary point of a broad spectrum of practical optimization problems such as the maximization of the minimum/sum rate, GEE and minimum EE, when the set of the feasible RIS coefficients adheres to convexity. In summary, the key conclusions of this work are:
@@ V Summary and Conclusions
  • The benefits of RISs increase when the packet lengths are reduced and/or the tolerable bit error rate is lower. The packet length can be related to the latency constraint, and the tolerable bit error rate represents the reliability constraint. Thus, these results show that RISs can be even more beneficial in URLLC systems than in non-URLLC systems.
@@
In future research, it would be interesting to develop advanced interference-management techniques such as RSMA and NOMA for MU-MIMO URLLC BCs. Moreover, studying the performance of other concepts/technologies for RIS, such as holographic RIS, active RIS, and BD-RIS, can be another promising direction for extending this work.

### v2

@@ IV-B2 Impact of the reliability constraint
Fig. 10 shows the average max-min EE versus $\epsilon$ for $P=10$ dB, $N_{{BS}}=4$, $N_{u}=4$, $K=2$, $L=1$, $M=1$, $n_{t}=256$ bits, and $N_{{RIS}}=20$. In this example, the RIS provides a significant gain, which increases with $\epsilon$. Again, RIS decreases the max-min EE when its elements are random. Moreover, we can observe that single-stream data transmission is suboptimal in this $4\times 4$ MIMO system. Indeed, the multi-stream systems communicating without RIS outperforms the RIS-aided single-stream data transmission.
@@ IV-B3 Impact of the packet length
Fig. 13 shows the average EE improvement by RISs versus $n_{t}$ for $P=10$ dB, $N_{{BS}}=2$, $N_{u}=2$, $K=2$, $L=2$, $M=2$, $\epsilon=10^{-5}$, and $N_{{RIS}}=20$. The average improvements reduce as $n_{t}$ increases. This indicates that the lower the tolerable latency, the higher gain the RIS can provide. In other words, RIS-aided systems become more energy efficient when a low latency is required, as in control channels, for example.
V Summary, Conclusions, and Future Research
An optimization framework was proposed for MU-MIMO RIS-aided systems with FBL by considering the NA for the rate expressions. To this end, we first calculated closed-form expressions for the FBL rate and then obtained suitable concave lower bounds for the FBL rates. Our proposed framework can be adapted to a large variety of MU-MIMO systems in which interference is treated as noise. Moreover, the framework can obtain a stationary point of a broad spectrum of practical optimization problems such as the maximization of the minimum/sum rate, GEE and minimum EE, when the set of the feasible RIS coefficients adheres to convexity. In summary, the key conclusions of this work are:
@@ V Summary, Conclusions, and Future Research
  • The benefits of RISs increase when the packet length is reduced and/or the tolerable bit error rate is lower. The packet length can be related to the latency constraint, and the tolerable bit error rate represents the reliability constraint. Thus, these results show that RISs can be even more beneficial in URLLC systems than in non-URLLC systems.
@@
In future research, it would be interesting to integrate advanced interference-management techniques, such as RSMA and NOMA, into MU-MIMO URLLC systems. In this regard, the solutions in, e.g., [36, 48], in combinations with the FBL rate expressions in Lemma 1 might be helpful. Another challenging research direction is to extend the optimization framework proposed in this work to scenarios with statistical or imperfect CSI. To this end, the robust designs in [12, 64] can be harnessed. Moreover, studying the performance of other concepts/technologies for RIS, such as holographic RIS [65], active RIS [66], BD-RIS [67, 68], and globally-passive RIS [69, 70] can be another promising direction for extending this work. Furthermore, considering RIS-aided cell-free URLLC systems is worth exploring in future research. Finally, another interesting line of research is to develop computationally more efficient resource allocation schemes that do not compromise performance.